solve sqrt(9x+55)=x+5
Answer by dabanfield(803)

solve sqrt(9x+55)=x+5
If we square both sides we have:
9x+55 = (x+5)^2
9x + 55 = x^2 + 10x + 25
x^2 + x - 30 = 0
(x+6)*(x-5) = 0
The solutions are when x+6 = 0 and/or x-5 = 0. That is x=-6 and/or x=5.